Calculate the volume in L of a rectangular solid with the following three length dimensions: L= 4.5 x 103 km, W = 2 x 10-4 Tm, H = 5.5 X 104 mm
1
Expert's answer
2020-03-24T14:01:27-0400

Some important conversions which I will be using here are:

  1. "1\ km=10^3 m"
  2. "1\ Tm=10^3 m"
  3. "1\ mm=10^{-2}m"
  4. "1\ m^3=10^3 Litres"

Volume="L\times B\times H=4.5\times10^3\times 10^3\times2\times10^{-4}\times10^3\times5.5\times10^{4}\times10^{-2}=49.5\times10^7m^3"

"=49.5\times10^7\times10^3 Litres=49.5\times10^{10} Litres"